OpenStudy (anonymous):

Can someone explain Im simplifying a slope intercept form equation. The original one is 3x - 2y = 6 The book says the answer is Y = 3/2x - 3 I get the minus 3 but why is it not negative 3/2

OpenStudy (anonymous):

is the book wrong?

OpenStudy (anonymous):

book is right, u try once again

OpenStudy (anonymous):

Your answer is not very helpful. If i knew why i would of done it along time ago. How is -3/-2 equal to a postive 3/2

OpenStudy (anonymous):

ok in a mint

OpenStudy (anonymous):

-3/-2= [-1*3]/[-1*2] =[-1/-1]*[3/2]=3/2 as [-1/-1] both numerator and denominator are the same they result in +1

OpenStudy (anonymous):

fine?

OpenStudy (anonymous):

Can i think about it like this? -3 in the numerator is the same as -3/1 and the -2 in the denominator is the same as -2/1 so we devide those to fractions my multiplying it by its reciporcal to get postive 3/2

OpenStudy (anonymous):

yes, but that is a little bit lenghty

OpenStudy (anonymous):

I dont understand in your answer where you are getting -1/-1

OpenStudy (anonymous):

u use twiddla?

OpenStudy (anonymous):

i dont know what that is

OpenStudy (anonymous):

look -3=-1*3

OpenStudy (anonymous):

i got that, yes

OpenStudy (anonymous):

-2=-1*2

OpenStudy (anonymous):

-3/-2=[-1*3]/[-1*2]

OpenStudy (anonymous):

=[-1/-1]*[3/2]

OpenStudy (anonymous):

so the -1,s cacnle and your left with 2/3

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

thank you!